40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 18 18:33:57 2014, ericq, Update, Green Locking, Yarm Green PDH yLoop.pdfcrap.pdf
    Reply  Tue Aug 19 01:01:36 2014, Jenne, Rana, Update, Green Locking, Yarm Green PDH YPDH_OLG.pdfYPDH_noise.pdf
       Reply  Tue Aug 19 18:32:40 2014, ericq, Update, Green Locking, Yarm Green PDH 
       Reply  Tue Aug 19 23:11:15 2014, Jenne, Update, Green Locking, Yarm Green PDH ErrAndCtrlSpectra_VoltsPerRtHz.png
       Reply  Wed Aug 20 02:38:41 2014, Jenne, Update, Green Locking, Yarm Green PDH - requirement NoiseConsideration.pdf
          Reply  Wed Aug 20 18:05:18 2014, Jenne, Update, Green Locking, Yarm Green PDH - requirement YPDH_noise.pdf
       Reply  Wed Aug 20 04:09:21 2014, ericq, Update, Green Locking, Xarm Green PDH Xbodes.pdfXspectra.pdfscopeSweep.jpg
          Reply  Wed Aug 20 16:10:43 2014, ericq, Update, Green Locking, Xarm Green PDH XspectraCombined.pdf
             Reply  Wed Aug 20 21:09:16 2014, ericq, Update, Green Locking, Xarm Green PDH scopeSweep.jpgXspectra.pdfYspectra.pdfXYcomp.pdf
Message ID: 10413     Entry time: Wed Aug 20 04:09:21 2014     In reply to: 10408     Reply to this: 10415
Author: ericq 
Type: Update 
Category: Green Locking 
Subject: Xarm Green PDH 

I've made a whole bunch of measurements on the Xarm green situation.

TL;DRs:

  • GTRX was around 0.55 for all of the measurements tonight. 
  • Based on where I saw gain peaking in the CLG, it looked like UGF was 1-2kHz. I cranked the gain to 10kHz, ~20dB gain peaking followed, making it hard to measure. Currently sitting at 5kHz-ish. 
  • Measured CLG with AG4395A, calibrated for injection point response, inferred OLG. 
  • Took various PSDs, still need to calibrate into physically meaningful units. 

Reasonable amounts of time were spent bending the AG4395 to my will; i.e. figuring out the calibration things Jenne and Rana did, finding the right excitation amplitude and profile that would leave the light steadily locked, and finding the right GPIB incantation for getting spectra in PSD units instead of power units. I'm nearing completion of a newer version of AG4395 scripts that have proper units, and pseudo-log spectra (i.e. logarithmically spaced linear sweeps)


Transfer functions

Here is too many traces on one plot showing parts of the OLTF for the x green PDH. One notable omission is the PD response (note to self:check model and bandwidth). The servo oddly seems to have a notch around 100k. My calibration for the CLG injection may not have been perfect, instead of flattening out at 0dB, I had 2dB residual. I tried to correct for it after the fact, assuming that certain regions were truly flat at 0dB, but I want to revisit it to be thorough. I found some old measurements of the Innolight PZT PM response, which claims to be in rad/V, and have included that on the plot. 

Xbodes.pdf

In the end, the mixer and PZT response make it look like getting over 10kHz bandwidth may be tough. Even finding a good higher modulation frequency to be able to scoot the LP up would leave us with the sharp slope in the PZT phase loss, and could cause bad gain peaking. Maybe it's worth thinking about a faster way of modulating the green light?


Noise Spectra

Tomorrow morning, I'll calibrate all the noise spectra I have into real units. These include:

  • In loop error signal and control signal spectra
  • Mixer output spectrum when PD is dark, and when mixer input is terminated
  • Servo out spectrum when PD is dark, and when servo input is terminated

However, looking at the floors, it occurs to me that I may have left the attenuation on the input too high, in an effort to protect the input the PDH box, which rails all the time when not locked to a 00 mode, sometimes even with the input terminated or open. It's kind of a pain that the agilent makes it really hard to see the data when you're in V/rtHz mode, because I should've caught this while measuring :/

I used a scope to capture a pdh signal happening, which will let me transform the mixer output into cavity motion. The control signal goes to the innolight PZT with a ~1MHz/V factor. Here are the uncalibrated plots, for now. 

scopeSweep.jpg

Xspectra.pdf

 

ELOG V3.1.3-